# A bound-preserving and conservative enriched Galerkin method for elliptic problems

### v1

@@ 6.3 Smooth solution
i.e., we have $a=0$ and $b=1$. In the following, we investigate the convergence of the method on a sequence of (nested) unstructured triangulations $\mathcal{T}_{h}$. As suggested by the theory, we choose $\beta=4$ (for the first two test cases) and the tolerance of the outer loop is $tol_{m}=10^{-12}$. Further, following [BGPV24], we choose the damping parameter $\omega=0.5$ and set $\alpha=1$ in (10).
@@
In Table 2, we present the results for the convergence of the method for different tolerances of the inner loop and the stabilization factor $\gamma=10$. We observe that the method converges for a wide range of tolerances of the inner loop, and the number of (outer) iterations $it$ is only mildly influenced by the choice of $tol_{n}$. Further, we see that the estimated order of convergence $eoc$ of the errors is optimal, i.e., we observe a quadratic rate for the error $\|u-u_{h}^{+}\|_{0,\Omega}$, and a linear rate for $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$. Note, that the $H^{1}$-error only considers the linear part in $V_{h}^{1}$ but not the constant part in $V_{h}^{0}$, i.e. $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}=\|\nabla(u-(u_{h}^{+})^{1})\|_{0,\Omega}$.
@@
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ $eoc$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$$\#its$
$496$ $5.33\cdot 10^{-3}$ (--- -) $3.44\cdot 10^{-1}$ (--- -) $9$
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) $8$
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) $6$
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) $2$
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) $1$
@@
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ $eoc$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$$\#its$
$496$ $5.33\cdot 10^{-3}$ (--- -) $3.44\cdot 10^{-1}$ (--- -) $5$
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) $5$
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) $4$
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) $2$
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) $1$
@@
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ $eoc$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$$\#its$
$496$ $5.33\cdot 10^{-3}$ (--- -) $3.44\cdot 10^{-1}$ (--- -) $3$
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) $2$
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) $2$
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) $2$
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) $1$
@@
In Figure 1 we present the number of iterations for different stabilization factors $\gamma$ and a tolerance of the inner loop set to $tol_{n}=10^{-9}$. In the left plot, the numbers are given for $\epsilon=10^{-5}$ and on the right for $\epsilon=1$. We observe that the number of iterations is barely influenced by the choice of $\gamma$, even for a small $\epsilon$. To motivate this behavior, we recall the stabilizing term in the bilinear form $a_{h}(\cdot,\cdot)$ given by $\langle\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\rangle$. We see that even in the case of small $\epsilon$ and a coarse mesh, since $\mu=1$ and $\beta\geq 4$, we get a relatively big penalization of the jumps of the solution even for a moderate $\gamma$. This would not be the case (particularly for vanishing $\epsilon$) if we use the stabilization $\gamma\frac{\epsilon}{h_{F}^{\beta}}$ instead, see also Remark 1, point 5.
$10^{3}$$10^{4}$$10^{5}$111122223333$|\mathcal{T}_{h}|$#$\#its,\epsilon=10^{-5}$$\gamma=10$$\gamma=10^{3}$$\gamma=10^{6}$$10^{3}$$10^{4}$$10^{5}$111122223333$|\mathcal{T}_{h}|$#$\#its,\epsilon=1$
Figure 1: Number of iterations for the example of Section 6.3 for different stabilization factors $\gamma$ and with $tol_{n}=10^{-9}$ and $tol_{m}=10^{-12}$.
@@
Although the theory requires that $\beta\geq 4$, we discuss in the following the convergence for different (smaller) choices $\beta=2,3,4$. We choose $tol_{n}=10^{-9}$ and set $\gamma=10$. Furthermore, in contrast to the previous example, we choose $\varepsilon=10^{-3}$, as our computations showed faster pre-asymptotic convergence of some errors for smaller choices. Consequently, we increased the diffusion coefficient to make the jumps more pronounced. In Table 3 we again present the error $\|\nabla(u-u_{h}^{+})\|_{L^{2}(\Omega)}$ but further present the values of the jump norm $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}=\|[\![(u_{h}^{+})^{0}]\!]\|_{\mathcal{F}_{h}}$ and the $L^{2}$-norm $\|(u_{h}^{+})^{0}\|_{L^{2}(\Omega)}$. We observe that the error $\|\nabla(u-u_{h}^{+})\|_{L^{2}(\Omega)}$ is independent of the choice of $\beta$ and converges with the optimal order. Note that the same conclusions can be made for the error $\|(u-u_{h}^{+})\|_{L^{2}(\Omega)}$, which is omitted here for brevity. The jump norm $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ shows a faster convergence. This suggests that the convergence analysis presented in Corollary 1 may be sharpened. More precisely, the numerical results indicate that $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ converges at a rate of order $h^{\beta}$. As in the previous paragraph, this observation can be motivated by the fact that for $\beta=1$, the stabilization essentially reduces to a standard interior penalty method, for which linear convergence of order $\mathcal{O}(h)$ is typically expected.
Since standard error estimates are often based on ellipticity with respect to $a_{h}$, as used in the first step of the proof of Lemma 9, we can argue that for $\beta>1$
@@
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$ $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ $eoc$ $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ $eoc$
$496$ $3.12\cdot 10^{-1}$ (--- -) $3.16\cdot 10^{-4}$ (--- -) $3.49\cdot 10^{-4}$ (--- -)
@@
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$ $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ $eoc$ $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ $eoc$
$496$ $3.12\cdot 10^{-1}$ (--- -) $4.03\cdot 10^{-5}$ (--- -) $3.23\cdot 10^{-5}$ (--- -)
@@
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$ $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ $eoc$ $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ $eoc$
$496$ $3.12\cdot 10^{-1}$ (--- -) $3.47\cdot 10^{-6}$ (--- -) $2.88\cdot 10^{-6}$ (--- -)
@@
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ $eoc$ $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ $eoc$ $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ $eoc$
$496$ $3.12\cdot 10^{-1}$ (--- -) $2.87\cdot 10^{-7}$ (--- -) $2.60\cdot 10^{-7}$ (--- -)
@@
Table 3: Error convergence for the example of Section 6.3 with the diffusion coefficient $\varepsilon=10^{-3}$, stabilization $\gamma=10$ and tolerances $tol_{n}=10^{-9}$ and $tol_{m}=10^{-12}$ and varying $\beta=1,2,3,4$.
@@ 6.4 Interior layer example
$f=\begin{cases}0&\text{ in }[\tfrac{1}{4},\tfrac{3}{4}]^{2},\\ 1&\text{ else}.\end{cases}$
Due to the discontinuous right-hand side and the small parameter $\varepsilon$ we expect that the solution will exhibit an interior layer.
To investigate this in more detail, Figure 2 shows the solution of the standard EG method and the solution $u_{h}^{+}$ of our proposed method for different stabilization parameters. Herein, we show results obtained using $\gamma=10$, but larger values of $\gamma$ give very similar results. For the standard method, we directly solved the problem (i.e., without using an iterative scheme) and employed $\beta=1$, $\gamma=10$, and $\alpha=0$ (no stabilization by (10)). The computations were made on a mesh with $|{\mathcal{T}_{h}}|=242$ and the same tolerances and damping parameters as in the previous section. We observe that the solution of the standard EG method exhibits a highly oscillatory behavior in the interior layer and fails to preserve the limits. In contrast, the solution of our proposed method is bound-preserving. We want to emphasize that, although the solution appears to be approximated solely by the linear Lagrange Finite element - similarly to [BGPV24] - our method is locally conservative since the piecewise constant part of $u_{h}^{+}$ is very small thanks to the over-penalization.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811111.21.21.21.2
Figure 2: Solutions of the interior layer example of Section 6.4 for the standard EG method (left) and the proposed method with $\gamma=10$ (right). For better visualization, the solution was cut along the line $[0,1]\times\{0.5\}$.
@@ 7 Conclusion
We have proposed a bound-preserving EG method whose solution is locally (and globally) conservative. A fundamental tool to achieve this is the way we have built the limiting process, by leaving the piecewise constant part free, and limiting the piecewise linear part in such a way that the sum respects the bounds given by the continuous problem. As a result, the stabilization needed to compensate for the ”unconstrained” part of the solution does not depend on the piecewise constant part of it, which implies conservation. In addition, it is important to insist on the fact that, although for the analysis a variational inequality was used, the method itself is not equivalent to a variational inequality (unlike [BGPV24, ABP24, BPT25]). The method is proven to to approximate smooth solutions with optimal convergence rates if the analytical solution respects pre-defined upper and lower bounds. If the solution is not smooth, our approach still respects the upper and lower bounds, is mass conservative, and convergent, but the significance of the penalty parameter $\gamma$ increases: a larger $\gamma$ suppresses oscillations stronger than a smaller $\gamma$. However, the nonlinear problem remains solvable independent of our tuning parameters $\gamma$ and $\beta$ (which we fix to $4$). In addition, the use of the splitting algorithm presented in Section 6.1 allows us to completely bypass the ill-conditioning of the linear systems that would arise were (9) to be solved in a monolithic way.
Thus, we have provided an efficient numerical framework to approximate diffusion–reaction problems. Naturally, some open problems should be tackled in future research. Some of these challenges include:

### v2

@@ 6.3 Smooth solution
i.e., we have $a=0$ and $b=1$. In the following, we investigate the convergence of the method on a sequence of (nested) unstructured triangulations $\mathcal{T}_{h}$. As suggested by the theory, we choose $\beta=4$ (for the first two test cases) and the tolerance of the outer loop is $tol_{m}=10^{-12}$. Further, following [BarrenecheaGPV23], we choose the damping parameter $\omega=0.5$ and set $\alpha=1$ in (10).
@@
In Table 2, we present the results for the convergence of the method for different tolerances of the inner loop and the penalty factor $\gamma=10$. We observe that the method converges for a wide range of tolerances of the inner loop, and the number of (outer) iterations $it$ is only mildly influenced by the choice of $tol_{n}$. Further, we see that the estimated order of convergence eoceoc of the errors is optimal, i.e., we observe a quadratic rate for the error $\|u-u_{h}^{+}\|_{0,\Omega}$, and a linear rate for $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$. Note, that the $H^{1}$-error only considers the linear part in $V_{h}^{1}$ but not the constant part in $V_{h}^{0}$, i.e. $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}=\|\nabla(u-(u_{h}^{+})^{1})\|_{0,\Omega}$.
@@
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ eoceoc $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ eoceoc #its\#its
$496$ $5.33\cdot 10^{-3}$ (--) $3.44\cdot 10^{-1}$ (--) 99
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) 88
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) 66
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) 22
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) 11
@@
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ eoceoc $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ eoceoc #its\#its
$496$ $5.33\cdot 10^{-3}$ (--) $3.44\cdot 10^{-1}$ (--) 55
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) 55
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) 44
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) 22
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) 11
@@
$|{\mathcal{T}_{h}}|$ $\|u-u_{h}^{+}\|_{0,\Omega}$ eoceoc $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ eoceoc #its\#its
$496$ $5.33\cdot 10^{-3}$ (--) $3.44\cdot 10^{-1}$ (--) 33
$1984$ $1.30\cdot 10^{-3}$ ($2.03$) $1.65\cdot 10^{-1}$ ($1.06$) 22
$7936$ $3.26\cdot 10^{-4}$ ($2.00$) $7.81\cdot 10^{-2}$ ($1.08$) 22
$31744$ $8.37\cdot 10^{-5}$ ($1.96$) $3.76\cdot 10^{-2}$ ($1.05$) 22
$126976$ $2.13\cdot 10^{-5}$ ($1.97$) $1.86\cdot 10^{-2}$ ($1.01$) 11
@@
In Figure 2 we present the number of iterations for different penalty factors $\gamma$ and a tolerance of the inner loop set to $tol_{n}=10^{-9}$. In the left plot, the numbers are given for $\epsilon=10^{-5}$ and on the right for $\epsilon=1$. We observe that the number of iterations is barely influenced by the choice of $\gamma$, even for a small $\epsilon$. To motivate this behavior, we recall the stabilizing term in the bilinear form $a_{h}(\cdot,\cdot)$ given by $\langle\gamma\frac{\epsilon+\mu h_{F}^{2}}{h_{F}^{\beta}}[\![w_{h}]\!],[\![v_{h}]\!]\rangle$. We see that even in the case of small $\epsilon$ and a coarse mesh, since $\mu=1$ and $\beta\geq 4$, we get a relatively big penalization of the jumps of the solution even for a moderate $\gamma$. This would not be the case (particularly for vanishing $\epsilon$) if we use the penalty $\gamma\frac{\epsilon}{h_{F}^{\beta}}$ instead, see also Remark 2.1, point 5.
[…]
Figure 2: Number of iterations for the example of Section 6.3 for different penalty factors $\gamma$ and with $tol_{n}=10^{-9}$ and $tol_{m}=10^{-12}$.
[…]
Figure 3: Number of iterations for the example of Section 6.3 for different factors $\beta$ and with $tol_{n}=10^{-9}$ and $tol_{m}=10^{-12}$.
@@
Although the theory requires that $\beta\geq 4$, we discuss in the following the convergence for different (smaller) choices $\beta=2,3,4$. We choose $tol_{n}=10^{-9}$ and set $\gamma=10$. Furthermore, in contrast to the previous example, we choose $\varepsilon=10^{-3}$, as our computations showed faster pre-asymptotic convergence of some errors for smaller choices. Consequently, we increased the diffusion coefficient to make the jumps more pronounced. In Table 3 we again present the error $\|\nabla(u-u_{h}^{+})\|_{L^{2}(\Omega)}$ but further present the values of the jump norm $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}=\|[\![(u_{h}^{+})^{0}]\!]\|_{\mathcal{F}_{h}}$ and the $L^{2}$-norm $\|(u_{h}^{+})^{0}\|_{L^{2}(\Omega)}$. We observe that the error $\|\nabla(u-u_{h}^{+})\|_{L^{2}(\Omega)}$ is independent of the choice of $\beta$ and converges with the optimal order. Note that the same conclusions can be made for the error $\|(u-u_{h}^{+})\|_{L^{2}(\Omega)}$, which is omitted here for brevity. The jump norm $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ shows a faster convergence. This suggests that the convergence analysis presented in Corollary 5.1 may be sharpened. More precisely, the numerical results indicate that $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ converges at a rate of order $h^{\beta}$. As in the previous paragraph, this observation can be motivated by the fact that for $\beta=1$, the penalty essentially reduces to a standard interior penalty method, for which linear convergence of order $\mathcal{O}(h)$ is typically expected.
Since standard error estimates are often based on ellipticity with respect to $a_{h}$, as used in the first step of the proof of Lemma 4.16, we can argue that for $\beta>1$
@@
In accordance to the discussions from the previous paragraph (”The choice of $\gamma$”), Figure 3 reports the number of outer iterations as a function of $\beta$ for a fixed $\gamma=10$. We make the following observations. First, in the right plot for $\epsilon=1$, all choices of $\beta$ provided a decreasing number of iterations with respect to the number of elements. Further, we see that the number of iterations on a fixed mesh also decreases as $\beta$ increases. This is explained by the stronger jump penalization for larger $\beta$, which drives the piecewise-constant component $u_{h}^{0}$ toward zero more rapidly (see Table 3); hence the outer update (which corrects $u_{h}^{0}$) becomes smaller and fewer outer iterations are needed to meet the tolerance.
However, in the left plot for $\epsilon=10^{-5}$, we see that for $\beta=1,2$ the iteration count does not decrease with respect to the number of elements, while for $\beta=3,4$ it does. Here we can clearly see the influence of $\gamma$ which has to chosen big enough (and obviously in dependence of $\epsilon$). Fortunately, as motivated by the theory, the choice $\beta=4$ provides a robust number of iterations also for small $\epsilon$.
@@
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ eoceoc $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ eoceoc $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ eoceoc
$496$ $3.12\cdot 10^{-1}$ (--) $3.16\cdot 10^{-4}$ (--) $3.49\cdot 10^{-4}$ (--)
@@
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ eoceoc $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ eoceoc $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ eoceoc
$496$ $3.12\cdot 10^{-1}$ (--) $4.03\cdot 10^{-5}$ (--) $3.23\cdot 10^{-5}$ (--)
@@
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ eoceoc $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ eoceoc $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ eoceoc
$496$ $3.12\cdot 10^{-1}$ (--) $3.47\cdot 10^{-6}$ (--) $2.88\cdot 10^{-6}$ (--)
@@
$|{\mathcal{T}_{h}}|$ $\|\nabla(u-u_{h}^{+})\|_{0,\Omega}$ eoceoc $\|(u_{h}^{+})^{0}\|_{0,\Omega}$ eoceoc $\|[\![u_{h}^{+}]\!]\|_{\mathcal{F}_{h}}$ eoceoc
$496$ $3.12\cdot 10^{-1}$ (--) $2.87\cdot 10^{-7}$ (--) $2.60\cdot 10^{-7}$ (--)
@@
Table 3: Error convergence for the example of Section 6.3 with the diffusion coefficient $\varepsilon=10^{-3}$, penalty $\gamma=10$ and tolerances $tol_{n}=10^{-9}$ and $tol_{m}=10^{-12}$ and varying $\beta=1,2,3,4$.
@@ 6.4 Interior layer example
$f=\begin{cases}0&\text{ in }[\tfrac{1}{4},\tfrac{3}{4}]^{2},\\ 1&\text{ else},\end{cases}$
while we impose homogeneous Dirichlet boundary conditions on the whole boundary $\partial\Omega$. The exact solution of this problem is not known analytically. However, we know that the solution satisfies $a=0\leq u\leq 1=b$ in $\Omega$. Due to the discontinuous right-hand side and the small parameter $\varepsilon$ we expect that the solution will exhibit an interior layer.
To investigate this in more detail, Figure 4 shows the solution of the standard EG method and the solution $u_{h}^{+}$ of our proposed method for different penalty parameters. Herein, we show results obtained using $\gamma=10$, but larger values of $\gamma$ give very similar results. For the standard method, we directly solved the problem (i.e., without using an iterative scheme) and employed $\beta=1$, $\gamma=10$, and $\alpha=0$ (no stabilization by (10)). The computations were made on a mesh with $|{\mathcal{T}_{h}}|=242$ and the same tolerances and damping parameters as in the previous section. We observe that the solution of the standard EG method exhibits a highly oscillatory behavior in the interior layer and fails to preserve the limits. In fact, the values of the solution on the left range from approximately 0.2-0.2 to $1.2$, as can be seen in the left panel of Figure 4. In contrast, the solution of our proposed method is bound-preserving. We want to emphasize that, although the solution appears to be approximated solely by the linear Lagrange Finite element - similarly to [BarrenecheaGPV23] - our method is locally conservative since the piecewise constant part of $u_{h}^{+}$ is very small thanks to the over-penalization.
One final point is that, since the constant part of $u_{h}^{+}$ is very small and the linear part strictly satisfying the homogeneous Dirichlet boundary conditions, the overall solution $u_{h}^{+}$ is very close to the boundary datum on the boundary. In contrast to that, the piecewise constants in the standard EG method (left panel) are not small conveying the impression that the standard EG method does not respect the Dirichlet boundary condition, as can be expected since it is only enforced via Nitsche’s method (guaranteeing that the boundary condition holds strictly in the limit $h\to 0$).
0.2-0.200.20.20.40.40.60.60.80.8111.21.2
Figure 4: Solutions of the interior layer example of Section 6.4 for the standard EG method (left) and the proposed method with $\gamma=10$ (right). For better visualization, the solution on the partial domain $0,1]\times[0.5,1]$ is shown here.
@@ 7 Conclusion
We have proposed a bound-preserving EG method whose solution is locally (and globally) conservative. A fundamental tool to achieve this is the way we have built the limiting process, by leaving the piecewise constant part free, and limiting the piecewise linear part in such a way that the sum respects the bounds given by the continuous problem. As a result, the stabilization needed to compensate for the “unconstrained” part of the solution does not depend on the piecewise constant part of it, which implies conservation. In addition, it is important to insist on the fact that, although for the analysis a variational inequality was used, the method itself is not equivalent to a variational inequality (unlike [BarrenecheaGPV23, AmiriBP23, BPT25]). The method is proven to approximate smooth solutions with optimal convergence rates if the analytical solution respects pre-defined upper and lower bounds. If the solution is not smooth, our approach still respects the upper and lower bounds, is mass conservative, and convergent, but the significance of the penalty parameter $\gamma$ increases: a larger $\gamma$ suppresses oscillations stronger than a smaller $\gamma$. However, the nonlinear problem remains solvable independent of our tuning parameters $\gamma$ and $\beta$ (which we fix to $4$). In addition, the use of the splitting algorithm presented in Section 6.1 allows us to completely bypass the ill-conditioning of the linear systems that would arise were (9) to be solved in a monolithic way.
Notably, our analysis requires $\beta\geq 2$ for the existence of discrete solutions and $\beta\geq 4$ for optimal convergence rates. However, numerical experiments indicate that the method performs well even for $\beta=1$, which is commonly used in interior penalty methods. This suggests that the theoretical requirements on $\beta$ may be too restrictive, and further investigation into this aspect could be a valuable direction for future research. In addition, the following problems are also open at the moment, and investigation on them is ongoing:
@@
  • The proposal of, possibly a posteriori, strategies to fix the stopping criteria for the nested fixed-point iterations in an adaptive way, with the aim of reducing computational complexity.